Two motorist start toward each other at 4:30 P.M. from towns 255 miles apart. If their respective average speeds are 40 and 45 ml/hr, at what time will they meet?
Please help me solve this problem.

Answer by mananth(16946)

Motorist I speed = 40 mph
Motorist II = 45 mph
They are travelling towards each other. effective speed = 40+45 =85 mph
Distance = speed * time
Time = distance /. speed
distance = 225 miles
speed = 85 mph
time= 225/85
= 2.65 hours
0.65 hours = 0.65*60 = 39 minutes
They started at 4.30 pm
add 2 hours 39 miutes to the starting time.
they meet at 7:09 pm